OpenStudy (anonymous):

Density of lead is 11.3kg m^-3.How much is the density of lead in g cm^-3 ? A 1.13x10^6 B 1.13x10^8 C 1.13x10^4 D 1.13x10^10 E 1.13x10^9 Just to double-check whether my answer is correct @ganeshie8

rvc (rvc):

1kg =? gm ?

rvc (rvc):

1kg = 10^3 gm correct?

rvc (rvc):

1m=10^2cm correct?

OpenStudy (anonymous):

correct

rvc (rvc):

so what is your ans?

OpenStudy (anonymous):

I say C

rvc (rvc):

\[1~kg\cdot~m^{-3}=10^3 \cdot~(10^2)^{-3}~gm~cm^{-3}\]

rvc (rvc):

i guess the options are wrong please check the options again

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

hold on

rvc (rvc):

please calculate the above calculation which i mentioned how much do you get?

rvc (rvc):

\(\rm 1~kg\cdot~m^{-3}=10^3 \cdot~(10^2)^{-3}~gm~cm^{-3}\)

OpenStudy (anonymous):

i get 1.13x10^-2

rvc (rvc):

correct so the answer is 0.0113 gm cm^-3

OpenStudy (anonymous):

i thought centi=10^-2

rvc (rvc):

1m=100cm

OpenStudy (anonymous):

Thank you @rvc :) i guess the options are wrong

rvc (rvc):

my pleasure yes the options are wrong :)